OpenStudy (anonymous):

Obtaining general solution: a.) ylnx(dx/dy)=[(y+1)/x]^2 b.) dy/dx=(xy+3x-y-3)/(xy-2x+4y-8) Need help. Thanks :)

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)

ganeshie8 (ganeshie8):

looks like separation of variables, did u try ?

ganeshie8 (ganeshie8):

\[\large y \ln x \dfrac{dx}{dy}=\left(\dfrac{y+1}{x}\right)^2\]

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

Ok. I'll try. Thanks! :)

OpenStudy (anonymous):

xlnxdx=[(y+1)^2dy]/y Is this correct??

ganeshie8 (ganeshie8):

\[\large x^2\ln x dx=\dfrac{(y+1)^2}{y} dy\]

ganeshie8 (ganeshie8):

yes integrate both sides

ganeshie8 (ganeshie8):

\[\large \int x^2\ln x dx=\int \dfrac{(y+1)^2}{y} dy\]

OpenStudy (anonymous):

Thank you!!! :)

ganeshie8 (ganeshie8):

np :) you need to work the left hand side by parts... hope you can finish it off

ganeshie8 (ganeshie8):

for the second problem, factor the right hand side first : http://www.wolframalpha.com/input/?i=factor+%28xy%2B3x-y-3%29%2F%28xy-2x%2B4y-8%29

ganeshie8 (ganeshie8):

good luck !

OpenStudy (anonymous):

Did I got it right? I obtained this equation.. 1/3(x^2)(lnx-1/3)=C

OpenStudy (anonymous):

I mean from the left side only..

ganeshie8 (ganeshie8):

\[\large \int x^2\ln x dx=\int \dfrac{(y+1)^2}{y} dy \] \[\large \frac{1}{3}x^3\ln x - \frac{x^3}{9}=\int \dfrac{y^2 + 2y + 1}{y} dy \] \[\large \frac{1}{3}x^3\ln x - \frac{x^3}{9}=\int y + 2 + \dfrac{ 1}{y} dy \] \[\large \frac{1}{3}x^3\ln x - \frac{x^3}{9}= \dfrac{y^2}{2} + 2y + \ln y + C \]

ganeshie8 (ganeshie8):

for the left side you should get : 1/3(x^3)(lnx-1/3)